情報集積設計学 2 単位 (選択)

Integrated Information System Design  福見稔 · 教授/ システム創生工学専攻知能情報システム工学コース基礎情報工学講座

島本隆 · 教授/ システム創生工学専攻電気電子創生工学コース知能電子回路講座

【授業目的】この講義の目的は VLSI の最新の設計手法を習得することです．【授業概要】高速情報処理のアルゴリズムを集積回路で実現するために，論理設

計，システム設計，大規模システム構成法およびシステム LSI の設計と製造法に関して講述する．とくに，ニューラル · ネットワークなどの並列 · 分散型情報処理の集積回路設計に関しても講義する．さらに， CAD を適用して VLSI を設計する方法に関して講義する． (赤松則男名誉教授) 高速アルゴリズムの構成法および集積回路設計を講義し，大規模システム LSI の構成と設計法を講義し，その応用に関しての講義も担当する． (福見稔教授) 並列 · 分散型情報処理の代表例としてニューラル · ネットワークの構成と集積回路設計法を講義し，その応用に関しての講義も担当する． (島本隆教授) ニューラル · ネットワークや遺伝的アルゴリズムなどを適用した最適化高速アルゴリズムの構成と高密度集積回路 (VLSI) の設計法を講義し，さらに， CAD の応用に関しての講義を担当する．

【授業形式】講義および演習【キーワード】 [キーワード] 【先行科目】 [先行科目] 【関連科目】 [関連科目] 【履修要件】博士前期課程で集積回路工学の単位を取得している必要がある．【履修上の注意】この講義を受講するには，最新の C-MOS 集積回路の知識を学ん

でいることが必要です．【到達目標】この講義は工学者と科学者が VLSI ニューロ · コンピューティングの

基礎的な知識を習得することを目標とする．【授業計画】

1. 組み込みソフトウエアのアーキテクチュア

2. リアル · タイムスケヂューリング手法

3. システム記述言語

4. 特定用途の集積回路

5. VLSI の消費電力と速度

6. 共有メモリと通信手段

7. キャッシュ · メモリと主記憶メモリ

8. システムのモデリングとドキュメンテーション

9. 集積回路システムの分割と性能 10. データ · フローグラフと有限状態機械

11. 動作記述言語と Spec C

12. 制御のデータ · フローグラフと機能合成

13. EEPROM 型プログラム可能シナップスを用いるニューロ · コンピューチングダイアグラム

14. レイアウト · パターンの例

15. ゲート · サイジングとタイミング · ドリブン

16. バウンダリ · スキャン概略とディレイ見積もり

【成績評価基準】試験と集積回路の設計で単位の認定を行う．

【教科書】システム LSI 設計， STARC 編集【参考書】赤松則男著; エレクトロニクス回路【WEB 頁】 http://www.A3.is.tokushima-u.ac.jp 【授業コンテンツ】 http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=216700 【対象学生】他学科，他学部学生も履修可能【連絡先】

⇒ 福見 (D 棟 210, 088-656-7510, fukumi@is.tokushima-u.ac.jp) ���� (オフィスアワー: 原則として，水曜日 15 時 ∼ 18 時，ただし年度により異なる場合があるので講義の際に指定する． )

【備考】講義と設計のスケジュールは個別に相談して下さい．

Integrated Information System Design 2 units (selection)

Minoru Fukumi · Professor / Information Science, Information Science and Intelligent Systems, Systems Innovation Engineering Takashi Shimamoto · Professor / Intelligent Networks and Computer Science, Electrical and Electronic Engineering, Systems Innovation Engineering

Target⟩ The aim of this lecture is to master the modern design technologies of

very large scale integrated circuits. Outline⟩ Very large scale integrated circuit (V-LSI) design and production method.

Using CAD technology, VLSI logic design, testing and fabrication are explained. Design of high-speed algorithm and parallel distributed processing system. Neural network and genetic algorithm for integrated circuit design.

Style⟩ Lecture and excercise �������⟩ [キーワード] Fundamental Lecture⟩ [先行科目] Relational Lecture⟩ [関連科目] Requirement⟩ It is necessary to get the unit of the mos integrated circuits in

master cource. Notice⟩ In order to get the unit of this lecture, the grduate cource students should

have learned the-state-of-the-art of the modern hardware technology, especially C-MOS integrated circuits.

Goal⟩ This lecture is desighned to provide engineers and scientists with an

introduction to the fieled of VLSI neurocomputing.

Schedule⟩

1. Embedded software architecture

2. Real-time schedure method

3. System description language

4. Application specific integrated circuits

5. Power consumption and speed of very large scale integrated circuits

6. Shared memory and communication mrthod

7. Cash memory and main memory

8. System modeling and documentation

9. Partitioning and performance

10. deta flow graph and finite state machine

11. Behavior description language and Spec C

12. Control deta-flow graph and function synthesis

13. Neural computing board diagram using EEPROM-style programmable synapses

14. Layout patern example 15. Gate-sizing wiring and timing driven

16. Boundary scan and delay estimation Evaluation Criteria⟩ Unit evaluation contains test and design of VLSI Textbook⟩ Hardware Annealing in Analog VLSI Neurocomputing, Kluer Academic

Publishers Reference⟩ Electronics Circuits, written by Norio Akamatsu

Webpage⟩ http://www.A3.is.tokushima-u.ac.jp

Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=216700

Student⟩ Able to be taken by student of other department and faculty

Contact⟩

⇒ Fukumi (D210, +81-88-656-7510, fukumi@is.tokushima-u.ac.jp) ���� (Office Hour: 原則として，水曜日 15 時 ∼ 18 時，ただし年度により異なる場合があるので講義の際に指定する． )

Note⟩ Lecturer will show the schedule of this lecture and design technologies.